Fama’s Llamas has a weighted average cost of capital of 9.5 percent. The company’s cost of equity is 14 percent, and its pretax cost of debt is 6.1 percent. The tax rate is 25 percent. What is the company’s target debt-equity ratio?

(Do not round intermediate calculations and round your answer to 4 decimal places, e.g., .1616.)
Answer #1

rate positively .. let me know if you need any clarification..

Lets assume the weight of equity = X. Therefore weight of debt = (1-x)
we will have below equation
X*14%+(1-X)*6.1%*(1-25%)= 9.50%
14%X+4.575%-4.575%X= 9.50%
9.425%X= 4.925%
X= 52.25%
equity weight = 52.25%
debt weight = 47.75%
debt/equity =                                  0.91
therefore answer =                                  0.91